7. (12 points) A time-domain equalizer is often used in digital communication to reduce the interferences between transmittedplease help me solve this problem as completely, and thoroughly, as possible. Thank you!
Answer #1

X(Z).7-1 21 sol xen] o 2 1 V, 2xtz) X(2) { 2.2 M2) y En] hox(z) 412) - hox(7) thi 77x(z) + h2 z xcz) Y(z) & hoth, zt the Z& [n] - S[nl] – S[n-2] + 2 S [n-3] & x[n-2] = 8 [n-2] - S En 3] +25[n- 47. putting & En] x [n ], a[n-2] in y En] we get y [n]11 oly. for 4[4.5 ML] - 6 -6 + Sto1( 2 be -b th.) to to 5 - 2 honhitha 5 - 2 + 1 - (-1) + ha o in s 5 = 2 + 1 tha 225-3 Th222Feel free to ask for any doubt you get.

Do rate the answer.